# Retrieval and Distill: A Temporal Data Shift-Free Paradigm for Online Recommendation System

### v4

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. INRODUCTION
@@
Click-through rate (CTR) prediction models play a significant role in modern recommendation systems. To enhance the accuracy of CTR models, deep learning-based approaches have been extensively employed (Covington et al., 2016; Shan et al., 2016; Wang et al., 2017, 2021). However, the dynamic environment faced by the recommendation systems, with the data distribution evolving over time, poses significant challenges to the effectiveness of these models (Zhu et al., 2023).
@@
To tackle this challenge, prior research has undertaken investigations from two distinct perspectives, namely incremental learning and behavior sequence modeling. On one hand, incremental learning techniques delve into updating the models that make these predictions as new data comes in, without having to start from scratch every time (Ye et al., 2022; Peng et al., 2021; Katsileros et al., 2022). This means the models learn and improve gradually as they get access to more recent data. However, the process of learning neural network parameters depends on iterative updates based on gradients, which slowly infuse supervision data into the model weights using a minimal learning rate. This poses a significant challenge for large parametric recommendation models in swiftly adapting to shifts in distribution. On the other hand, behavior sequence modeling focuses on understanding how users’ behaviors change over time by looking at the sequence of actions they take. In this area, various network structures have been proposed, including recurrent neural networks (RNNs) (Hidasi et al., 2015, 2016), convolutional neural networks (CNNs) (Tang and Wang, 2018; Yuan et al., 2019), and memory networks (Pi et al., 2019; Ren et al., 2019). Moreover, the attention mechanism has emerged as the predominant approach for modeling item dependencies, with several notable contributions such as SASRec (Kang and McAuley, 2018), DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), and BERT4Rec (Sun et al., 2019). These studies aim to identify sequential patterns in user behavior, thereby facilitating the prediction of future preferences based on historical activities. However, these studies fail to directly tackle the issue of distribution shift during training time, and these high-capacity models significantly increase the online workload and inference burden.
@@
Figure 1. The phenomenon of temporal data shift on a real dataset collected from Tmall. The test set consists of data after Nov. 10, 2015. The x-axis represents the date from which we add all data up to Nov. 10 to train a CTR model, and the y-axis represents the model’s performance on the test set. We found that the model’s performance peaks after adding just a few final days of data. However, further adding data from earlier periods leads to a decline in the model’s performence on the test set due to the inconsistency in data distribution between this earlier data and the test set.
@@
All of the above-mentioned models either update the model using the latest data (Guo et al., 2017; Wang et al., 2017) or use past shifting data as a reference for retrieval (Qin et al., 2020; Qi et al., 2020), without directly using the shifting data for training model parameters. However, directly using the temporally shifting data for model training is challenging, as an example illustrated in Figure 1 based on a public dataset of Tmall, an international e-commerce platform. In this dataset, data spanning Nov. 11-12, 2015 is employed as the test set. The x-axis represents the training set from the indicated date to Nov. 10, 2015. We train a common deep CTR model PNN (Qu et al., 2018) using different training sets to make predictions. The y-axis indicates the AUC performance of the model on the test set. We can observe that utilizing data over Nov. 1-10, 2015 for predictions on the test set yields the best results, whereas adding data from earlier time periods will negatively impact the results due to data shift. Therefore, it is crucial to find an effective method that can directly utilize these data for model training.
@@
Inspired by the recent proposed retrieval-based tabular data prediction models (Qin et al., 2021; Zheng et al., 2023), we found that similar data from historical records can significantly enhance the predictive capability and generalizability. In other words, although the probability distribution modeled by the CTR is time-variant, the probability distribution of the association between current user behavior data and similar historical data can remain invariant over time if the search space is fixed. If we design a neural network to model the probability distribution of data similar to the current query, this neural network is temporal data shift-free. Thus, the shifting data can be used to train such a neural network. By incorporating this neural network with the original CTR model, we can obtain an aggregated prediction model that has been trained with shifting data.
@@
In this paper, we propose the Retrieval and Distill paradigm (RAD) to address the challenge of ineffective training of a CTR model over shifting data. The RAD paradigm generally consists of two major components, namely the Retrieval Framework and the Distill Framework. We first establish a Retrieval Framework to assist the original model in making predictions using similar data within the shifting data. We refer to the neural network in Retrieval Framework that retrieves and outputs the logits of similar data as the relevance network. By pre-training this relevance network using shifting data and then merging the original CTR model and relevance network, we can significantly enhance the predictive capability of the original CTR model. However, the added relevance network in the Retrieval Framework, due to its retrieval process (such as BM25 (Robertson et al., 1995)), is difficult to deploy online due to its inference time, similar to other retrieval-based CTR models (Qin et al., 2021; Zheng et al., 2023). To address this difficulty, we further build the Distill Framework to distill knowledge from the relevance network using shifting data, obtaining a parameterized student network called search-distill module. By aggregating the original CTR model and the search-distill module, we finetune the Distill Framework with unshifting data, e.g., the latest training data. This yields the final model that can be deployed online for light workload and efficient inference.
@@
2. PRELIMINARIES
In this section, we formulate the problem and introduce the notations. In the context of recommendation systems, we represent the dataset as a tabular structure, denoted by $\mathcal{D}$, comprising $F$ feature columns $\mathcal{C}=\{C_{i}\}_{i=1}^{F}$ alongside a singular label column $\mathcal{Y}$. The feature columns encapsulate attributes such as ”Age,” ”City,” and ”Category.” Each individual entry within table $\mathcal{D}$ corresponds to a sample $s_{z}$. A sample is characterized by an amalgamation of multiple features along with a label, articulated as $s_{z}=(x_{z},y_{z})$, where $x_{z}=\{c_{i}^{z}\}_{i=1}^{F}$. Accordingly, table $\mathcal{D}$ is defined as an ensemble of $N$ samples, expressed as $\mathcal{D}=\{s_{z}\}_{z=1}^{N}$.
@@ 2. PRELIMINARIES
(3) $\mathcal{L}_{\text{pretrain}}=\sum_{x_{z}\in\mathcal{D}_{\text{shifting}}}\text{CE}(y^{z},\hat{y}_{T}^{z}),$
@@
(4) $\mathcal{L}_{\text{retrieval}}=\sum_{x_{z}\in\mathcal{D}_{\text{train}}}\text{CE}(y^{z},\hat{y}_{R}^{z}),$
@@
(8) $\mathcal{L}_{\text{KD}}=\sum_{x_{z}\in\mathcal{D}_{\text{shifting}}}\text{MSE}(r(x),R(x))=\sum_{x_{z}\in\mathcal{D}_{\text{shifting}}}\text{MSE}(w_{r}^{z},w_{R}^{z}),$
@@
(9) $\mathcal{L}_{\text{distill}}=\sum_{x_{z}\in\mathcal{D}_{\text{train}}}\text{CE}(y^{z},\hat{y}_{D}^{z}),$
@@ 3. The RAD Paradigm
Figure 2. This figure illustrates the original model and its combination with the two components of the RAD paradigm. In CTR tasks, the original model typically takes tabular data as input and outputs a probability indicating the likelihood of a click. The RAD paradigm initially employs a retrieval framework to assist the original model in making predictions using shifting data. However, considering the performance requirements of the CTR model, we designed the RAD’s distill framework to eliminate the search process in the retrieval framework, thereby meeting the online inference time requirements.
Figure 3. Retrieval framework.
Figure 4. Distill Process.
Figure 5. Distill Framework.
3.1. Overview
Traditional CTR models have not accounted for the data shift issue, and they cannot enhance performance by merely adding shifting data (Juan et al., 2016; Naumov et al., 2019; Gu et al., 2020). Training only with unshifting data neglects a significant amount of information. Our proposed Retrieval and Distill paradigm aims to devise a universal solution that utilizes shifting data to enhance the performance of existing models. RAD is generally divided into two parts. The first part incorporates a Retrieval Framework to introduce shifting data into the datastore during training. Concurrently, as per theorem 2.1, we pretrain the model’s relevance network with shifting data and then transfer it for use. This approach substantially improves the model’s performance without altering its original structure. However, including a non-parametric search component in the relevance network poses deployment challenges for online recommendation systems. Therefore, the second part is a Distill Framework tasked with transforming the relevance network into a parametric neural network model. We distill a simple neural network (search-distill module) using shifting data and the relevance network as the teacher network. Ultimately, we merge the search-distill module with the original model and finetune it using unshifting data to yield the final online deployable model. The following sections will elucidate these modules in detail.
3.2. Retreival Framework
@@
3.2.1. Framework Model Structure
@@
Relevance Network. The relevance network features an embedding layer shared with input $x_{z}$, a search engine, and an attention-based aggregation layer. Within our framework, we utilize shifting tabular data entries as the search space, with BM25 (Robertson et al., 1995) serving as our search algorithm.
@@
(11) $\text{RankScore}(x_{z},x_{d})$ $=\sum_{j=1}^{M}\text{IDF}(a_{j}^{z})\cdot\mathbf{1}(x_{j}^{z}=x_{j}^{d}),$
(12) $\text{IDF}(x_{j}^{z})$ $=\log\frac{N-N(a_{j}^{z})+0.5}{N(a_{j}^{z})+0.5},$
@@
(14) $w_{R}=\sum_{k=1}^{K}\alpha_{k}\cdot x_{k},\quad s_{k}\in\mathcal{S}(x_{z}),$
@@
Teacher Network. The teacher network is a Retrieval Framework that does not include the original model. From Figure 3, we can observe that the teacher network forms the right portion of the diagram. The teacher network serves as an intermediary bridge for training the relevance network. The teacher network takes $x_{z}$ and the relevance network output as inputs, producing a probability as output. Drawing inspiration from RIM (Qin et al., 2021), we incorporate a factorization machine module within the aggregation layer of the teacher network. This aggregation layer is designed to effectively merge the input $x_{z}$ with the aggregated $w_{R}$, thus improving the teacher network’s capability to discover the hidden relationships between $x_{z}$ and the returned $S(x_{z})$ from the search engine.
3.2.2. Pretrain with Shifting Data
@@
3.2.3. Finetune with Unshifting Data
@@
3.3. Distill Framework
Although the Retrieval Framework alone allows us to significantly enhance the performance of various existing CTR models by utilizing shifting data, this improvement depends on the relevance network’s inclusion of BM25 search, which can be computationally intensive due to the typically large volume of $\mathcal{D}_{\text{shifting}}$ data and the often linear time complexity of the search process. This presents a substantial challenge for deploying the Retrieval Framework in an online environment. To optimize the time complexity of the Retrieval Framework, we propose a Distill Framework designed to distill the relevance network and remove its search component. The resulting module is termed the search-distill module.
Next, we start by delineating the structure of our model, followed by a description of the Distill Framework’s training process. Within the Retrieval Framework’s pretraining step, we achieve a pretrained relevance network. In the Distill Framework, this network undergoes distillation using $\mathcal{D}_{\text{shifting}}$. Subsequently, the distilled search-distill module is combined with the original model through an integrated multilayer neural network for aggregation and subsequent finetuning using $\mathcal{D}_{training}$.
3.3.1. Framework Model Structure
@@
3.3.2. Pretrain and Distill with Shifting Data
@@
3.3.3. Finetune with Unshifting Data
@@ 4. Experiments
4.1. Datasets
In order to evaluate the performance of RAD, we conduct extensive experiments for CTR prediction tasks on three real-world large-scale recommendation datasets from Alibaba and Ant Group, i.e., Taobao, Tmall and Alipay [^1] https://tianchi.aliyun.com/dataset/x, where ‘x’ is ‘649’, ‘42’, and ‘53’ for Taobao, Tmall, and Alipay, respectively.. These datasets consist of sequential user-item interactions in e-commerce scenes. Table  1 is the information for our split dataset.
Datasets Fields Split # Instances Start time End time
Taobao 4 $\mathcal{D}_{\text{shifting}}$ 145416702 2017-11-25 2017-12-01
$\mathcal{D}_{\text{train}}$ 2754415 2017-12-02 2017-12-02
$\mathcal{D}_{\text{test}}$ 2734177 2017-12-03 2017-12-03
Tmall 9 $\mathcal{D}_{\text{shifting}}$ 68646472 2015-05-11 2015-10-31
$\mathcal{D}_{\text{train}}$ 6679610 2015-11-01 2015-11-10
$\mathcal{D}_{\text{test}}$ 7055119 2015-11-11 2015-11-12
Alipay 6 $\mathcal{D}_{\text{shifting}}$ 82090360 2015-07-01 2015-10-31
$\mathcal{D}_{\text{train}}$ 264843 2015-11-21 2015-11-24
$\mathcal{D}_{\text{test}}$ 431646 2015-11-25 2015-11-30
Table 1. Datasets
4.2. Evaluation Metrics
@@
4.3. Compared Methods
To evaluate the effectiveness of our proposed architecture, we choose six commonly used and strong models as backbones, which can be grouped into two classes.
4.3.1. Feature interaction based Models
The first class includes some traditional feature interaction based tabular models. DNN represents multilayer perceptron (MLP), a simple feedforward artificial neural network. DeepFM (Guo et al., 2017) is an architecture that combines the factorization machine and deep neural network. PNN (Qu et al., 2018) is a product-based neural network which captures both inter-field interactive patterns and high-order feature interactions.
4.3.2. Sequential based Models
The second class contains several sequential end-to-end deep models. Transformer (Vaswani et al., 2017) is a sequence-to-sequence model based on self-attention mechanism. Here, we use its encoder to model users’ history behaviors. DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019)are two attention-based sequential models, while DIEN adds a recurrent neural network based on DIN.

### v5

@@ Abstract
1 INRODUCTION
@@
Click-through rate (CTR) prediction models play a significant role in modern recommendation systems. To enhance the accuracy of CTR models, deep learning-based approaches have been extensively employed [3][28][33][34]. However, the dynamic environment faced by the recommendation systems, with the data distribution evolving over time, poses significant challenges to the effectiveness of these models [47].
@@
To tackle this challenge, prior research has undertaken investigations from two distinct perspectives, namely incremental learning and behavior sequence modeling. On one hand, incremental learning techniques delve into updating the models that make these predictions as new data comes in, without having to start from scratch every time [40][19][15]. This means the models learn and improve gradually as they get access to more recent data. However, the process of learning neural network parameters depends on iterative updates based on gradients, which slowly infuse supervision data into the model weights using a minimal learning rate. This poses a significant challenge for large parametric recommendation models in swiftly adapting to shifts in distribution. On the other hand, behavior sequence modeling focuses on understanding how users’ behaviors change over time by looking at the sequence of actions they take. In this area, various network structures have been proposed, including recurrent neural networks (RNNs) [9][10], convolutional neural networks (CNNs) [31][41], and memory networks [21][26]. Moreover, the attention mechanism has emerged as the predominant approach for modeling item dependencies, with several notable contributions such as SASRec [14], DIN [46], DIEN [45], and BERT4Rec [29]. These studies aim to identify sequential patterns in user behavior, thereby facilitating the prediction of future preferences based on historical activities. However, these studies fail to directly tackle the issue of distribution shift during training time, and these high-capacity models significantly increase the online workload and inference burden.
@@
Figure 1: The phenomenon of temporal data shift on a real dataset collected from Tmall. The test set consists of data after Nov. 10, 2015. The x-axis represents the date from which we add all data up to Nov. 10 to train a CTR model, and the y-axis represents the model’s performance on the test set. We observed that model performance peaked when trained on temporally adjacent data, with performance degrading as the training period extended further from the test set timeframe.
@@
All of the above-mentioned models either update the model using the latest data [6, 33] or use past shifting data as a reference for retrieval [23, 22], without directly using the shifting data for training model parameters. However, directly using the temporally shifting data for model training is challenging, as an example illustrated in Figure 1 based on a public dataset of Tmall, an international e-commerce platform. In this dataset, data spanning Nov. 11-12, 2015 is employed as the test set. The x-axis represents the training set from the indicated date to Nov. 10, 2015. We train a common deep CTR model PNN [25] using different training sets to make predictions. The y-axis indicates the AUC performance of the model on the test set. We can observe that utilizing data over Nov. 1-10, 2015 for predictions on the test set yields the best results, whereas adding data from earlier time periods will negatively impact the results due to data shift. Therefore, it is crucial to find an effective method that can directly utilize these data for model training.
@@
Inspired by the recent proposed retrieval-based tabular data prediction models [24, 44], we found that similar data from historical records can significantly enhance the predictive capability and generalizability. In other words, although the probability distribution modeled by the CTR is time-variant, the probability distribution of the association between current user behavior data and similar historical data can remain invariant over time if the search space is fixed. If we design a neural network to model the probability distribution of data similar to the current query, this neural network is temporal data shift-free. Thus, the shifting data can be used to train such a neural network. By incorporating this neural network with the original CTR model, we can obtain an aggregated prediction model that has been trained with shifting data.
@@
In this paper, we propose the Retrieval and Distill paradigm (RAD) to address the challenge of ineffective training of a CTR model over shifting data. The RAD paradigm generally consists of two major components, namely the Retrieval Framework and the Distill Framework. We first establish a Retrieval Framework to assist the original model in making predictions using similar data within the shifting data. We refer to the neural network in Retrieval Framework that retrieves and outputs the logits of similar data as the relevance network. By pre-training this relevance network using shifting data and then merging the original CTR model and relevance network, we can significantly enhance the predictive capability of the original CTR model. However, the added relevance network in the Retrieval Framework, due to its retrieval process (such as BM25 [27]), is difficult to deploy online due to its inference time, similar to other retrieval-based CTR models [24][44]. To address this difficulty, we further build the Distill Framework to distill knowledge from the relevance network using shifting data, obtaining a parameterized student network called search-distill module. By aggregating the original CTR model and the search-distill module, we finetune the Distill Framework with unshifting data, e.g., the latest training data. This yields the final model that can be deployed online for light workload and efficient inference.
@@
2 PRELIMINARIES
In this section, we formulate the problem and introduce the notations. In the context of recommendation systems, we represent the dataset as a tabular structure, denoted by $\mathcal{D}$, comprising $F$ feature columns $\mathcal{C}=\{C_{i}\}_{i=1}^{F}$ alongside a singular label column $\mathcal{Y}$. The feature columns encapsulate attributes such as "Age," "City," and "Category." Each individual entry within table $\mathcal{D}$ corresponds to a sample $s_{z}$. A sample is characterized by an amalgamation of multiple features along with a label, articulated as $s_{z}=(x_{z},y_{z})$, where $x_{z}=\{c_{i}^{z}\}_{i=1}^{F}$. Accordingly, table $\mathcal{D}$ is defined as an ensemble of $N$ samples, expressed as $\mathcal{D}=\{s_{z}\}_{z=1}^{N}$.
@@ 2 PRELIMINARIES
$\mathcal{L}_{\text{pretrain}}=\sum_{x_{z}\in\mathcal{D}_{\text{shifting}}}\text{CE}(y^{z},\hat{y}_{T}^{z}),$ (3)
@@
$\mathcal{L}_{\text{retrieval}}=\sum_{x_{z}\in\mathcal{D}_{\text{train}}}\text{CE}(y^{z},\hat{y}_{R}^{z}),$ (4)
@@
$\mathcal{L}_{\text{KD}}=\sum_{x_{z}\in\mathcal{D}_{\text{shifting}}}\text{MSE}(r(x),R(x))=\sum_{x_{z}\in\mathcal{D}_{\text{shifting}}}\text{MSE}(w_{r}^{z},w_{R}^{z}),$ (8)
@@
$\mathcal{L}_{\text{distill}}=\sum_{x_{z}\in\mathcal{D}_{\text{train}}}\text{CE}(y^{z},\hat{y}_{D}^{z}),$ (9)
@@ 3 The RAD Paradigm
Figure 2: This figure illustrates the original model and its combination with the two components of the RAD paradigm. In CTR tasks, the original model typically takes tabular data as input and outputs a probability indicating the likelihood of a click. The RAD paradigm initially employs a retrieval framework to assist the original model in making predictions using shifting data. However, considering the performance requirements of the CTR model, we designed the RAD’s distill framework to eliminate the search process in the retrieval framework, thereby meeting the online inference time requirements.
Figure 3: Retrieval framework.
Figure 4: Distill Process.
Figure 5: Distill Framework.
3.1 Overview
Traditional CTR models have not accounted for the data shift issue, and they cannot enhance performance by merely adding shifting data [12][18][5]. Training only with unshifting data neglects a significant amount of information. Our proposed Retrieval and Distill paradigm aims to devise a universal solution that utilizes shifting data to enhance the performance of existing models. RAD is generally divided into two parts. The first part incorporates a Retrieval Framework to introduce shifting data into the datastore during training. Concurrently, as per theorem 2.1, we pretrain the model’s relevance network with shifting data and then transfer it for use. This approach substantially improves the model’s performance without altering its original structure. However, including a non-parametric search component in the relevance network poses deployment challenges for online recommendation systems. Therefore, the second part is a Distill Framework tasked with transforming the relevance network into a parametric neural network model. We distill a simple neural network (search-distill module) using shifting data and the relevance network as the teacher network. Ultimately, we merge the search-distill module with the original model and finetune it using unshifting data to yield the final online deployable model. The following sections will elucidate these modules in detail.
3.2 Retreival Framework
@@
3.2.1 Framework Model Structure
@@
Relevance Network. The relevance network features an embedding layer shared with input $x_{z}$, a search engine, and an attention-based aggregation layer. Within our framework, we utilize shifting tabular data entries as the search space, with BM25 [27] serving as our search algorithm.
@@
$\text{RankScore}(x_{z},x_{d})$ $=\sum_{j=1}^{M}\text{IDF}(a_{j}^{z})\cdot\mathbf{1}(x_{j}^{z}=x_{j}^{d}),$ (11)
$\text{IDF}(x_{j}^{z})$ $=\log\frac{N-N(a_{j}^{z})+0.5}{N(a_{j}^{z})+0.5},$ (12)
@@
$w_{R}=\sum_{k=1}^{K}\alpha_{k}\cdot x_{k},\quad s_{k}\in\mathcal{S}(x_{z}),$ (14)
@@
Teacher Network. The teacher network is a Retrieval Framework that does not include the original model. From Figure 3, we can observe that the teacher network forms the right portion of the diagram. The teacher network serves as an intermediary bridge for training the relevance network. The teacher network takes $x_{z}$ and the relevance network output as inputs, producing a probability as output. Drawing inspiration from RIM [24], we incorporate a factorization machine module within the aggregation layer of the teacher network. This aggregation layer is designed to effectively merge the input $x_{z}$ with the aggregated $w_{R}$, thus improving the teacher network’s capability to discover the hidden relationships between $x_{z}$ and the returned $S(x_{z})$ from the search engine.
3.2.2 Pretrain with Shifting Data
@@
3.2.3 Finetune with Unshifting Data
@@
3.3 Distill Framework
To optimize the time complexity of the Retrieval Framework, we propose a Distill Framework that distills the relevance network (pretrained via the Retrieval Framework) into a search-distill module by removing its computationally intensive BM25 search component. This module is then integrated with the original model through a multilayer neural network, aggregated, and finetuned using $\mathcal{D}_{\text{training}}$, enabling efficient deployment in online environments.
3.3.1 Framework Model Structure
@@
3.3.2 Pretrain and Distill with Shifting Data
@@
3.3.3 Finetune with Unshifting Data
@@ 4 Experiments
4.1 Datasets
In order to evaluate the performance of RAD, we conduct extensive experiments for CTR prediction tasks on three real-world large-scale recommendation datasets from Alibaba and Ant Group, i.e., Taobao, Tmall and Alipay [^1] https://tianchi.aliyun.com/dataset/x, where ‘x’ is ‘649’, ‘42’, and ‘53’ for Taobao, Tmall, and Alipay, respectively.. These datasets consist of sequential user-item interactions in e-commerce scenes. The datasets contain timestamped user behaviors. We split them chronologically: the newest data for testing, the oldest for retrieval ($\mathcal{D}_{\text{shifting}}$), and the middle segment for training ($\mathcal{D}_{\text{training}}$). This mimics real-world scenarios where models must adapt to evolving preferences while leveraging historical data.
4.2 Evaluation Metrics
@@
4.3 Compared Methods
To evaluate our framework, we select six strong baseline models grouped into two categories: (1) Feature interaction models: DNN (MLP), DeepFM [6] (combines FM and DNN), and PNN [25] (product-based feature interaction); (2) Sequential models: Transformer encoder [32], DIN [46], and DIEN [45] (DIN with RNN), which model user behavior sequences via attention mechanisms.